# Conformal Constrained Policy Optimization for Cost-Effective LLM Agents

### v1

@@ 1 Introduction
Specifically, we consider an agentic setting, where a policy orchestrates a small LLM and a large LLM across a series of rounds to answer the question, with the goal of minimizing the expected cost subject to a user-specified reliability level. To formalize the reliability constraint, we use conformal prediction (shafer2008tutorial), which provides theoretical guarantees by modifying the system to output sets of labels (called prediction sets) instead of individual labels. Conformal prediction provides the coverage guarantee, which says the prediction set contains the true label with high probability. In our setting, we can only cover the true answer if some orchestration strategy does so; thus, our reliability constraint says the following holds with high probability: if some sequence of actions produces the correct answer, then our prediction set contains the true label.
Thus, the orchestration policy must select among LLM agents to satisfy the coverage guarantee while minimizing average cost (we implicitly constrain prediction set size). Prior work has addressed only parts of this goal. For instance, chain-of-thought prompting (wei2022chain), ReAct (yao2023react), debate protocols, and RL-based LLM agent frameworks (zhou2025sweet; bensal2025reflect) improve accuracy but ignore cost. Alternatively, FrugalGPT (chen2024frugalgpt) reduces API expenses but does not guarantee reliability. Finally, uncertainty-aware methods (ren2023robots; han2024towards) calibrate model confidences but rely on ad-hoc techniques to perform model selection and do not seek to minimize cost. No existing approach unifies these desiderata.
We propose Conformal Constrained Policy Optimization (CCPO), which combines recent advances in online conformal prediction (angelopoulos2024online) with constrained policy optimization (achiam2017constrained). Specifically, we seek to train a conformal policy that outputs a prediction set over actions to minimize cost subject to a conformal constraint; an optional penalty can also be imposed on prediction set size. However, training a set-based action policy is infeasible; instead, we convert the conformal policy into a stochastic one (i.e., uniform distribution over the action set) and train this policy instead. A key challenge is that the conformal policy visits a different state-action distribution compared to the stochastic policy; to correct for this shift, we use importance weighted V-trace targets and policy gradients (espeholt2018impala). Finally, we apply online conformal prediction (angelopoulos2024online) to scale the action sets to guarantee coverage.
We empirically evaluate our approach on the free-form HotpotQA (yang2018hotpotqa) and multiple-choice MMLU (hendrycks2020measuring) datasets, showing that CCPO reduces total cost by up to 30% while satisfying target coverage guarantees, outperforming state-of-the-art cost-minimization baselines.
@@ 2 Related Work
One way to improve the accuracy of LLM agents is to instruct them to perform reasoning. Chain-of-Thought (CoT) explicitly elicits step-by-step explanations (wei2022chain), and ReAct interleaves reasoning with tool calls (yao2023react). Extensions incorporate uncertainty quantification—e.g., UALA estimates predictive variance before committing to an answer (han2024towards). They also propose to reduce cost by making early-exit decisions in multi-turn chains (lu2025runaway) and role-based multi-agent editing pipelines (wan2025mamm).
Rather than using hand-written prompts, an alternative strategy is to train control policies. min2025self first trains a Process Reward Model to score each intermediate thought and then learns a binary tabular policy—“request help” vs. “proceed”. SWEET-RL (zhou2025sweet) tackles multi-turn tasks by using external feedback as contextual states, yielding better credit assignment. bensal2025reflect leverages a self-reflection mechanism for self-improvement. However, while these methods improve reliability or fluency, they do not provide any guarantees. In contrast, we aim to provide reliability guarantees using conformal prediction.
@@
There has been work on imposing safety constraints in reinforcement learning. TRPO constrains each update by a KL trust region (schulman2015trust); CPO augments this approach with an explicit cost bound for one-step feasibility (achiam2017constrained). Shielding methods block actions that violate specifications (alshiekh2018safe), while Lagrangian penalties trade reward and risk (tessler2019reward). bastani2021safe combines a learned policy with a backup policy, sacrificing reward when necessary to ensure safety.
A complementary line of work tackles stable learning from off-policy data. Retrace clips importance ratios for stable Q-learning from replay (munos2016safe); IMPALA’s V-trace extends this to distributed actor–learner systems (espeholt2018impala). In purely offline data, CQL (kumar2020conservative) and IQL (kostrikov2021offline) suppress over-optimistic estimates with regularization.
@@
Online conformal prediction extends classical, exchangeability‐based CP to sequential or adversarial data streams by tracking the conformity threshold, largely through online optimization. ACI (gibbs2021adaptive) introduces an adaptive quantile estimator that preserves marginal coverage when the underlying distribution evolves. MVP (bastani2022practical) achieves adversarially robust, group-conditional (“multivalid”) coverage efficiently with tight prediction sets. angelopoulos2024online propose a weight-decayed quantile tracker that gives online conformal prediction “best-of-both-worlds” guarantees—O(T)O($\sqrt${T}) coverage regret for adversarial streams and near-optimal set sizes for i.i.d. data.
@@
stutz2021learning proposes a method to simulate split-conformal prediction within each mini-batch and backpropagate through the procedure, enabling the training of optimal score functions for classification problems. A final batch calibration step restores the coverage guarantee, yielding tighter confidence sets than conventional post-hoc conformal prediction. However, optimizing the conformal score function remains unexplored in sequential settings such as ours.
@@ 4 Conformal CPO
To update $\pi$, we adapt the trust‐region method from CPO (achiam2017constrained), yielding consistent improvement in our objective while maintaining our coverage constraint. Taking into account the KL‐ball search in CPO and the contractive validity of V-trace (see Section 6), we optimize the stochastic conformal policy $S_{\pi,\kappa}$ as the target policy. Concretely, for any score function $\pi$ and threshold $\kappa$, we define
@@
To bridge the gap between $\pi$ (which is used to collect trajectories) and $S_{\pi,\kappa}$, we employ V‐trace off‐policy correction as in espeholt2018impala. At each iteration, rollouts under $\pi$ produce tuples $(o_{t},a_{t},r_{t},c_{t})$, where $r_{t}$ and $c_{t}$ are the reward and constraint values on step $t$, respectively. In our POMDP in Eq. (1), we have reward values
@@
Next, we describe our trust–region policy update. We let $S_{k}=S_{\pi_{k},\kappa_{k}}$ and $C_{k}=C_{\pi_{k},\kappa_{k}}$ denote the stochastic and conformal policies, respectively, on the $k$th iteration of optimization. Then, as in achiam2017constrained, we iteratively solve the following approximation of Eq. (1) (we describe how $\kappa$ is updated later):
@@
$J_{C}^{\pi,\kappa}\leq\Pr[Y^{*}\in C_{\pi,\kappa}(Q)]-\Pr[Y^{*}\in\mathcal{Y}(Q)].$
@@
$\approx\mathds{E}_{\pi}\left[\left(\prod_{t=1}^{T}\rho_{t}\cdot|C_{\pi,\kappa}(o_{t})|\right)\mathds{1}\{S_{\pi,\kappa}(Q)=Y^{*}\}\right]$
@@
$\bar{J}_{C}^{\pi,\kappa}$ $=\bar{J}^{\pi,\kappa}_{C,0}-\Pr[Y^{*}\in\mathcal{Y}(Q)]$
$\bar{J}^{\pi,\kappa}_{C,0}$ $=\mathds{E}_{\pi}\left[\left(\prod_{t=1}^{T}\rho_{t}\cdot|C_{\pi,\kappa}(o_{t})|\right)\mathds{1}\{S_{\pi,\kappa}(Q)=Y^{*}\}\right].$
@@
The threshold $\kappa$ is updated after each episode to provably maintain the coverage guarantee without relying on convergence of the policy. When solving for $\pi_{k+1}$ in Eq. (7), we fix $\kappa=\kappa_{k}$. Afterwards, we update $\kappa_{k}$ based on the coverage of $C_{\pi_{k+1},\kappa_{k}}$ using the algorithm in angelopoulos2024online, that $\kappa_{k+1}$ is
@@
Assuming $\sum_{t}\eta_{t}=\infty$ and $\sum_{t}\eta_{t}^{2}<\infty$, then angelopoulos2024online show that in the i.i.d. setting, the coverage converges to $1-\alpha$ even though $\pi$ is being updated in a streaming fashion, as long as it converges. If additionally $\eta_{t}\propto t^{-1/2-\xi}$ for some $\xi\in(0,1/2)$, then in the adversarial setting the long-run coverage error is bounded as $O(T^{-1/2+\xi})$ (angelopoulos2024online, Theorem 1). While we guarantee coverage for the final policy $\pi_{K}$ assuming $\pi_{k}$ converges, we perform a batch calibration on $\pi_{K}$ (i.e., use traditional conformal prediction (shafer2008tutorial) on a held-out calibration set to select a final value $\kappa_{K+1}$) to ensure coverage for $\pi_{K}$ even without this assumption.
@@ 5.1 Experimental Setup
We evaluate our approach on two challenging question‐answering benchmarks. HotpotQA (yang2018hotpotqa) is a commonsense QA dataset requiring multi‐hop reasoning and often multiple Wikipedia searches to answer each question. MMLU (hendrycks2020measuring) is a multiple‐choice QA dataset spanning academic and professional subjects such as biology, mathematics, and physics. Following UALA (han2024towards), we use Wikipedia search as a tool for HotpotQA. For MMLU, UALA uses the Google API as a tool. However, the cost of Google API is \$5 per 1,000 calls, which far exceeds our budget; thus, we apply CoT with deterministic decoding (temperature 0.0) in the first step and temperature 1.0 in the subsequent steps. For HotpotQA, we train on 1,000 examples, while for MMLU, we train on 560 examples. For both, we use 200 examples for batch calibration and 200 for testing.
@@
  • LLM-EXIT (lu2025runaway): This baseline uses a single LLM both to provide answers and to decide whether to proceed to the next round.  [^1] For exit policies, we report matched coverage by summing the percentages of correct and unsolvable questions, ensuring a fair comparison.
  • GPT-4o-guided UALA: In UALA (han2024towards), both uncertainty estimates and answers are generated by the same model, which tend to be biased and inaccurate for smaller models. To align with our approach, we adopt GPT-4o to quantify uncertainty. Then, this baseline follows the same steps as our approach, except instead of using a learned policy, it applies fixed thresholds on uncertainty to choose the action.
@@ Discussion.
First, cost-aware policy optimization efficiently minimizes cost while maintaining the desired coverage, achieving orders of magnitude cost reduction compared to LLM-based methods with similar coverage. The uncertainty-thresholding method with UALA is also effective, consistent with the findings of chen2024frugalgpt. Nevertheless, policy optimization methods still offer a clear performance advantage, as they explicitly incorporate cost information and can potentially leverage additional contextual features for even stronger decision-making.
@@ 6 Discussion
To guarantee convergence and a unique fixed point for the critic, we clip the importance weights so that the V-trace operator becomes a contraction mapping. Specifically, espeholt2018impala show that with $\gamma\bar{c}<1$, the V-trace operator $\mathcal{R}V$ is a contraction mapping in the sup-norm. Thus, setting $\bar{c}=\bar{\rho}=1$ preserves the contraction property and introduces only negligible bias in the critic update. For policy updates, we apply clipping only to the reward gradient to reduce variance, while leaving the cost gradient unclipped. Clipping $g_{c}$ would introduce bias, so we avoid it. Overall, this clipping scheme ensures that the critic remains stable and that the CPO gradients are both low-variance and nearly unbiased.
@@
Second, espeholt2018impala argue that for $\mathcal{R}V$ to be a contraction mapping, the policy must satisfy $\mathds{E}_{\mu}\left[\frac{\pi_{\text{new}}(a_{t}\mid o_{t})}{\pi_{\text{old}}(a_{t}\mid o_{t})}\right]=1$. Using a stochastic policy ensures the importance weights are bounded, thereby preserving convergence of the V-trace method.
@@ Acknowledgements
8 Prompt Templates
8.1 Guide Model Template
Evaluate the answer.
- If correct, reply "Yes".
- If incorrect, reply "No [correct answer]" (only the correct answer, nothing else).
Q: {question}
A: {answer}
Your response:
8.2 LLM EXIT Template
You will be given a question and its historical answers. You must decide whether to exit.
Instructions:
Evaluate the current answer history and determine whether it meets any of the following conditions:
1. The recent responses are repetitive, or no valid answer has been given.
2. The most recent answer is correct, and no further attempts are necessary.
Input:
Question: {question}
Previous Answers: {answers}
Output:
Exactly one of [EXIT] or [CONTINUE], and nothing else.
Answer:
8.3 LLM-as-a-Policy Template
System Prompt:
You are a decision-making assistant. When given:
• A question
• An answer from a base model
• A judgement from a GPT-4o model in the format ‘‘agreement (yes/no), gpt_answer’’
• An uncertainty score from the GPT-4o model
Your job is to pick one of these actions:
1. Use the base model’s answer.
2. Use the GPT-4o model’s answer.
3. Go to the next round (neither answer is acceptable).
Your goal is to maintain high decision quality and approximately 0.9 coverage over accepted answers.
Output only the action number (‘‘1’’, ‘‘2’’, or ‘‘3’’) and nothing else.
User Content:
Question: {question}
Base answer: {base_ans}
GPT-4o judgement (agreement, answer): {gpt_judgement}
GPT-4o uncertainty: {gpt_uncertainty}
Action:
9 Additional Results
Policy Cost (cents) Coverage Avg. Len. Set Size
Random 6.811 0.578 1.44 1.00
GPT-4o 19.25 0.810 1.16 1.00
LLaMA-2-7b 4.524 0.685 1.02 1.00
UALA 4.635 0.850 1.00 1.00
CPO 4.711 0.840 1.00 1.00
CPO batch 4.663 0.841 1.00 1.00
CPO online 4.709 0.819 1.00 1.00
CCPO ($\lambda=0$) 4.640 0.862 1.00 1.00
CCPO ($\lambda=1e^{-4}$) 4.647 0.815 1.00 1.00
Table 6: HotpotQA results with LLaMA-2-7B, $\alpha=0.2$.
Figure 6: HotpotQA coverage and cost with LLaMA-2-7B base model over 100 splits, $\alpha=0.2$.
Policy Cost (cents) Coverage Avg. Len. Set Size
Random 1.411 0.610 1.51 1.00
GPT-4o 18.77 0.845 1.12 1.00
LLaMA-3.2-3b 4.495 0.605 1.00 1.00
UALA 13.25 0.970 3.00 3.00
CPO 4.722 0.846 1.02 1.00
CPO batch 8.717 0.979 1.97 3.66
CPO online 10.02 0.894 2.11 3.68
CCPO ($\lambda=0$) 8.126 0.954 1.86 3.39
CCPO ($\lambda=1e^{-4}$) 8.575 0.951 1.95 3.20
Table 7: HotpotQA results with LLaMA-3.2-3b, $\alpha=0.05$.
Figure 7: HotpotQA coverage and cost with LLaMA-3.2-3b base model over 100 splits, $\alpha=0.05$.
Policy Cost (cents) Coverage Avg. Len. Set Size
Random 12.99 0.545 1.53 1.00
GPT-4o 27.15 0.887 1.31 1.00
LLaMA-2-7b 8.712 0.520 1.12 1.00
UALA 7.612 0.889 1.00 1.00
CPO 7.521 0.860 1.00 1.00
CPO batch 8.024 0.886 1.00 1.00
CPO online 7.643 0.828 1.00 0.95
CCPO ($\lambda=0$) 7.612 0.801 1.00 0.90
CCPO ($\lambda=2e^{-4}$) 7.670 0.805 1.00 0.90
Table 8: MMLU results with LLaMA-2-7B, $\alpha=0.2$.
Figure 8: MMLU coverage and cost with LLaMA-2-7B base model over 100 splits, $\alpha=0.2$.
Policy Cost (cents) Coverage Avg. Len. Set Size
Random 12.74 0.607 1.55 1.00
GPT-4o 26.87 0.913 1.19 1.00
LLaMA-3-2-3b 8.547 0.580 1.03 1.00
UALA 7.755 0.839 1.00 1.00
CPO 7.121 0.805 1.00 1.00
CPO batch 7.767 0.805 1.00 0.91
CPO online 7.620 0.818 1.00 0.98
CCPO ($\lambda=0$) 7.735 0.806 1.00 1.02
CCPO ($\lambda=2e^{-4}$) 7.742 0.806 1.00 0.89
Table 9: MMLU results with LLaMA-3.2-3b, $\alpha=0.2$.
Figure 9: MMLU coverage and cost with LLaMA-3.2-3b base model over 100 splits, $\alpha=0.2$.
10 Training Curves
Costs
Refer to caption Refer to caption Refer to caption Refer to caption
Coverage Surrogate Violation
Refer to caption Refer to caption Refer to caption Refer to caption
Training Coverage
(a) LLaMA-2-7b, $\alpha=0.2$
(b) LLaMA-2-7b, $\alpha=0.1$
(c) LLaMA-3.2-3b, $\alpha=0.1$
(d) LLaMA-3.2-3b, $\alpha=0.05$
Figure 10: Overall CCPO training performance on HotpotQA dataset, $\lambda=0$.
Coverage
Refer to caption Refer to caption Refer to caption Refer to caption
Average Length
(a) LLaMA-2-7b, $\alpha=0.2$
(b) LLaMA-2-7b, $\alpha=0.1$
(c) LLaMA-3.2-3b, $\alpha=0.1$
(d) LLaMA-3.2-3b, $\alpha=0.05$
Figure 11: CCPO validation performance on HotpotQA dataset, $\lambda=0$.
Train Coverage
Refer to caption Refer to caption Refer to caption Refer to caption
Validation Coverage
Refer to caption Refer to caption Refer to caption Refer to caption
Validation Length
(a) LLaMA-2-7b, $\alpha=0.2$
(b) LLaMA-2-7b, $\alpha=0.1$
(c) LLaMA-3.2-3b, $\alpha=0.1$
(d) LLaMA-3.2-3b, $\alpha=0.05$
Figure 12: CPO performance on HotpotQA dataset.
Costs
Refer to caption Refer to caption Refer to caption Refer to caption
Coverage Surrogate Violation
Refer to caption Refer to caption Refer to caption Refer to caption
Training Coverage
(a) LLaMA-2-7b, $\alpha=0.2$,
$\lambda=1e^{-4}$.
(b) LLaMA-2-7b, $\alpha=0.1$,
$\lambda=2e^{-4}$.
(c) LLaMA-3.2-3b, $\alpha=0.1$,
$\lambda=1e^{-4}$.
(d) LLaMA-3.2-3b, $\alpha=0.05$,
$\lambda=1e^{-4}$.
Figure 13: Overall CCPO training performance on HotpotQA dataset with varying $\lambda$.
Coverage
Refer to caption Refer to caption Refer to caption Refer to caption
Average Length
(a) LLaMA-2-7b, $\alpha=0.2$,
$\lambda=1e^{-4}$.
(b) LLaMA-2-7b, $\alpha=0.1$,
$\lambda=2e^{-4}$.
(c) LLaMA-3.2-3b, $\alpha=0.1$,
$\lambda=1e^{-4}$.
(d) LLaMA-3.2-3b, $\alpha=0.05$,
$\lambda=1e^{-4}$.
Figure 14: CCPO validation performance on HotpotQA dataset with varying $\lambda$.
Costs
Refer to caption Refer to caption Refer to caption Refer to caption
Coverage Surrogate Violation
Refer to caption Refer to caption Refer to caption Refer to caption
Training Coverage
(a) LLaMA-2-7b, $\alpha=0.2$
(b) LLaMA-2-7b, $\alpha=0.1$
(c) LLaMA-3.2-3b, $\alpha=0.2$
(d) LLaMA-3.2-3b, $\alpha=0.1$
Figure 15: Overall CCPO training performance on MMLU dataset, $\lambda=0$.
Coverage
Refer to caption Refer to caption Refer to caption Refer to caption
Average Length
(a) LLaMA-2-7b, $\alpha=0.2$
(b) LLaMA-2-7b, $\alpha=0.1$
(c) LLaMA-3.2-3b, $\alpha=0.2$
(d) LLaMA-3.2-3b, $\alpha=0.1$
Figure 16: CCPO validation performance on MMLU, $\lambda=0$.
Train Coverage
Refer to caption Refer to caption Refer to caption Refer to caption
Validation Coverage
Refer to caption Refer to caption Refer to caption Refer to caption
Validation Length
(a) LLaMA-2-7b, $\alpha=0.2$
(b) LLaMA-2-7b, $\alpha=0.1$
(c) LLaMA-3.2-3b, $\alpha=0.2$
(d) LLaMA-3.2-3b, $\alpha=0.1$
Figure 17: CPO performance on MMLU dataset.
Costs
Refer to caption Refer to caption Refer to caption Refer to caption
Coverage Surrogate Violation
Refer to caption Refer to caption Refer to caption Refer to caption
Training Coverage
(a) LLaMA-2-7b, $\alpha=0.2$,
$\lambda=2e^{-4}$.
(b) LLaMA-2-7b, $\alpha=0.1$,
$\lambda=1e^{-4}$.
(c) LLaMA-3.2-3b, $\alpha=0.2$,
$\lambda=1e^{-4}$.
(d) LLaMA-3.2-3b, $\alpha=0.1$,
$\lambda=1e^{-4}$.
Figure 18: Overall CCPO training performance on MMLU dataset with varying $\lambda$.
Coverage
Refer to caption Refer to caption Refer to caption Refer to caption
Average Length
(a) LLaMA-2-7b, $\alpha=0.2$,
$\lambda=2e^{-4}$.
(b) LLaMA-2-7b, $\alpha=0.1$,
$\lambda=1e^{-4}$.
(c) LLaMA-3.2-3b, $\alpha=0.2$,
$\lambda=1e^{-4}$.
(d) LLaMA-3.2-3b, $\alpha=0.1$,
$\lambda=1e^{-4}$.
Figure 19: CCPO validation performance on MMLU with varying $\lambda$.

### v2

@@ 1 Introduction
Specifically, we consider an agentic setting, where a policy orchestrates a small LLM and a large LLM across a series of rounds to answer the question, with the goal of minimizing the expected cost subject to a user-specified reliability level. To formalize the reliability constraint, we use conformal prediction (Shafer and Vovk 2008), which provides theoretical guarantees by modifying the system to output sets of labels (called prediction sets) instead of individual labels. Conformal prediction provides the coverage guarantee, which says the prediction set contains the true label with high probability. In our setting, we can only cover the true answer if some orchestration strategy does so; thus, our reliability constraint says the following holds with high probability: if some sequence of actions produces the correct answer, then our prediction set contains the true label.
Thus, the orchestration policy must select among LLM agents to satisfy the coverage guarantee while minimizing average cost (we implicitly constrain prediction set size). Prior work has addressed only parts of this goal. For instance, chain-of-thought prompting (Wei et al. 2022), ReAct (Yao et al. 2023), debate protocols, and RL-based LLM agent frameworks (Zhou et al. 2025; Bensal et al. 2025) improve accuracy but ignore cost. Alternatively, FrugalGPT (Chen et al. 2024) reduces API expenses but does not guarantee reliability. Finally, uncertainty-aware methods (Ren et al. 2023; Han et al. 2024) calibrate model confidences but rely on ad-hoc techniques to perform model selection and do not seek to minimize cost. No existing approach unifies these desiderata.
We propose Conformal Constrained Policy Optimization (CCPO), which combines recent advances in online conformal prediction (Angelopoulos et al. 2024) with constrained policy optimization (Achiam et al. 2017). Specifically, we seek to train a conformal policy that outputs a prediction set over actions to minimize cost subject to a conformal constraint; an optional penalty can also be imposed on prediction set size. However, training a set-based action policy is infeasible; instead, we convert the conformal policy into a stochastic one (i.e., uniform distribution over the action set) and train this policy instead. A key challenge is that the conformal policy visits a different state-action distribution compared to the stochastic policy; to correct for this shift, we use importance weighted V-trace targets and policy gradients (Espeholt et al. 2018). Finally, we apply online conformal prediction (Angelopoulos et al. 2024) to scale the action sets to guarantee coverage.
We empirically evaluate our approach on the free-form HotpotQA (Yang et al. 2018) and multiple-choice MMLU (Hendrycks et al. 2021) datasets, showing that CCPO reduces total cost by up to 30% while satisfying target coverage guarantees, outperforming state-of-the-art cost-minimization baselines.
@@ 2 Related Work
One way to improve the accuracy of LLM agents is to instruct them to perform reasoning. Chain-of-Thought (CoT) explicitly elicits step-by-step explanations (Wei et al. 2022), and ReAct interleaves reasoning with tool calls (Yao et al. 2023). Extensions incorporate uncertainty quantification—e.g., UALA estimates predictive variance before committing to an answer (Han et al. 2024). They also propose to reduce cost by making early-exit decisions in multi-turn chains (Lu et al. 2025) and role-based multi-agent editing pipelines (Wan et al. 2025).
Rather than using hand-written prompts, an alternative strategy is to train control policies. Min et al. (2025) first trains a Process Reward Model to score each intermediate thought and then learns a binary tabular policy—“request help” vs. “proceed”. SWEET-RL (Zhou et al. 2025) tackles multi-turn tasks by using external feedback as contextual states, yielding better credit assignment. Bensal et al. (2025) leverages a self-reflection mechanism for self-improvement. However, while these methods improve reliability or fluency, they do not provide any guarantees. In contrast, we aim to provide reliability guarantees using conformal prediction.
@@
There has been work on imposing safety constraints in reinforcement learning. TRPO constrains each update by a KL trust region (Schulman et al. 2015); CPO augments this approach with an explicit cost bound for one-step feasibility (Achiam et al. 2017). Shielding methods block actions that violate specifications (Alshiekh et al. 2018), while Lagrangian penalties trade reward and risk (Tessler et al. 2019). Bastani et al. (2021) combines a learned policy with a backup policy, sacrificing reward when necessary to ensure safety.
A complementary line of work tackles stable learning from off-policy data. Retrace clips importance ratios for stable Q-learning from replay (Munos et al. 2016); IMPALA’s V-trace extends this to distributed actor–learner systems (Espeholt et al. 2018). In purely offline data, CQL (Kumar et al. 2020) and IQL (Kostrikov et al. 2022) suppress over-optimistic estimates with regularization.
@@
Online conformal prediction extends classical, exchangeability‐based CP to sequential or adversarial data streams by tracking the conformity threshold, largely through online optimization. ACI (Gibbs and Candes 2021) introduces an adaptive quantile estimator that preserves marginal coverage when the underlying distribution evolves. MVP (Bastani et al. 2022) achieves adversarially robust, group-conditional (“multivalid”) coverage efficiently with tight prediction sets. Angelopoulos et al. (2024) propose a weight-decayed quantile tracker that gives online conformal prediction “best-of-both-worlds” guarantees—O(T)O($\sqrt${T}) coverage regret for adversarial streams and near-optimal set sizes for i.i.d. data.
@@
Stutz et al. (2022) proposes a method to simulate split-conformal prediction within each mini-batch and backpropagate through the procedure, enabling the training of optimal score functions for classification problems. A final batch calibration step restores the coverage guarantee, yielding tighter confidence sets than conventional post-hoc conformal prediction. However, optimizing the conformal score function remains unexplored in sequential settings such as ours.
@@ 4 Conformal CPO
To update $\pi$, we adapt the trust‐region method from CPO (Achiam et al. 2017), yielding consistent improvement in our objective while maintaining our coverage constraint. Taking into account the KL‐ball search in CPO and the contractive validity of V-trace (see Section 6), we optimize the stochastic conformal policy $S_{\pi,\kappa}$ as the target policy. Concretely, for any score function $\pi$ and threshold $\kappa$, we define
@@
To bridge the gap between $\pi$ (which is used to collect trajectories) and $S_{\pi,\kappa}$, we employ V‐trace off‐policy correction as in Espeholt et al. (2018). At each iteration, rollouts under $\pi$ produce tuples $(o_{t},a_{t},r_{t},c_{t})$, where $r_{t}$ and $c_{t}$ are the reward and constraint values on step $t$, respectively. In our POMDP in Eq. (1), we have reward values
@@
Next, we describe our trust–region policy update. We let $S_{k}=S_{\pi_{k},\kappa_{k}}$ and $C_{k}=C_{\pi_{k},\kappa_{k}}$ denote the stochastic and conformal policies, respectively, on the $k$th iteration of optimization. Then, as in Achiam et al. (2017), we iteratively solve the following approximation of Eq. (1) (we describe how $\kappa$ is updated later):
@@
$J_{C}^{\pi,\kappa}\leq\Pr[Y^{*}\in C_{\pi,\kappa}(Q)]+\Pr[Y^{*}\notin\mathcal{Y}(Q)].$
@@
$\approx\mathds{E}_{\pi}\left[\left(\prod_{t=1}^{T}\rho_{t}\cdot|C_{\pi,\kappa}(o_{t})|\right)\mathds{1}\{\pi(Q)=Y^{*}\}\right]$
@@
$\bar{J}_{C}^{\pi,\kappa}$ $=\bar{J}^{\pi,\kappa}_{C,0}+\Pr[Y^{*}\notin\mathcal{Y}(Q)]$
$\bar{J}^{\pi,\kappa}_{C,0}$ $=\mathds{E}_{\pi}\left[\left(\prod_{t=1}^{T}\rho_{t}\cdot|C_{\pi,\kappa}(o_{t})|\right)\mathds{1}\{\pi(Q)=Y^{*}\}\right].$
@@
The threshold $\kappa$ is updated after each episode to provably maintain the coverage guarantee without relying on convergence of the policy. When solving for $\pi_{k+1}$ in Eq. (7), we fix $\kappa=\kappa_{k}$. Afterwards, we update $\kappa_{k}$ based on the coverage of $C_{\pi_{k+1},\kappa_{k}}$ using the algorithm in Angelopoulos et al. (2024), that $\kappa_{k+1}$ is
@@
Assuming $\sum_{t}\eta_{t}=\infty$ and $\sum_{t}\eta_{t}^{2}<\infty$, then Angelopoulos et al. (2024, Theorem 4 and Corollary 2) show that in the i.i.d. setting, the coverage converges to $1-\alpha$ even though $\pi$ is being updated in a streaming fashion, as long as it converges. If additionally $\eta_{t}\propto t^{-1/2-\xi}$ for some $\xi\in(0,1/2)$, then in the adversarial setting the long-run coverage error is bounded as $O(T^{-1/2+\xi})$ (Angelopoulos et al. 2024, Theorem 1). While we guarantee coverage for the final policy $\pi_{K}$ assuming $\pi_{k}$ converges, we perform a batch calibration on $\pi_{K}$ (i.e., use traditional conformal prediction (Shafer and Vovk 2008) on a held-out calibration set to select a final value $\kappa_{K+1}$) to ensure coverage for $\pi_{K}$ even without this assumption.
@@ 5.1 Experimental Setup
We evaluate our approach on two challenging question‐answering benchmarks. HotpotQA (Yang et al. 2018) is a commonsense QA dataset requiring multi‐hop reasoning and often multiple Wikipedia searches to answer each question. MMLU (Hendrycks et al. 2021) is a multiple‐choice QA dataset spanning academic and professional subjects such as biology, mathematics, and physics. Following UALA (Han et al. 2024), we use Wikipedia search as a tool for HotpotQA. For MMLU, UALA uses the Google API as a tool. However, the cost of Google API is \$5 per 1,000 calls, which far exceeds our budget; thus, we apply CoT with deterministic decoding (temperature 0.0) in the first step and temperature 1.0 in the subsequent steps. For HotpotQA, we train on 1,000 examples, while for MMLU, we train on 560 examples. For both, we use 200 examples for batch calibration and 200 for testing.
@@
  • LLM-EXIT (Lu et al. 2025): This baseline uses a single LLM both to provide answers and to decide whether to proceed to the next round.  [^1] For exit policies, we report matched coverage by summing the percentages of correct and unsolvable questions, ensuring a fair comparison.
  • GPT-4o-guided UALA: In UALA (Han et al. 2024), both uncertainty estimates and answers are generated by the same model, which tend to be biased and inaccurate for smaller models. To align with our approach, we adopt GPT-4o to quantify uncertainty. Then, this baseline follows the same steps as our approach, except instead of using a learned policy, it applies fixed thresholds on uncertainty to choose the action.
@@ Discussion.
First, cost-aware policy optimization efficiently minimizes cost while maintaining the desired coverage, achieving orders of magnitude cost reduction compared to LLM-based methods with similar coverage. The uncertainty-thresholding method with UALA is also effective, consistent with the findings of Chen et al. (2024). Nevertheless, policy optimization methods still offer a clear performance advantage, as they explicitly incorporate cost information and can potentially leverage additional contextual features for even stronger decision-making.
@@ 6 Discussion
To guarantee convergence and a unique fixed point for the critic, we clip the importance weights so that the V-trace operator becomes a contraction mapping. Specifically, Espeholt et al. (2018, proof of Remark 3) show that with $\gamma\bar{c}<1$, the V-trace operator $\mathcal{R}V$ is a contraction mapping in the sup-norm. Thus, setting $\bar{c}=\bar{\rho}=1$ preserves the contraction property and introduces only negligible bias in the critic update. For policy updates, we apply clipping only to the reward gradient to reduce variance, while leaving the cost gradient unclipped. Clipping $g_{c}$ would introduce bias, so we avoid it. Overall, this clipping scheme ensures that the critic remains stable and that the CPO gradients are both low-variance and nearly unbiased.
@@
Second, Espeholt et al. (2018, proof of Remark 3) argue that for $\mathcal{R}V$ to be a contraction mapping, the policy must satisfy $\mathds{E}_{\mu}\left[\frac{\pi_{\text{new}}(a_{t}\mid o_{t})}{\pi_{\text{old}}(a_{t}\mid o_{t})}\right]=1$. Using a stochastic policy ensures the importance weights are bounded, thereby preserving convergence of the V-trace method.